OpenStudy (anonymous):

find the measures of 2 supplementary angles if one angle is 12 more than 3 times the other find the measures of 2 supplementary angles if one angle is 12 more than 3 times the other @Mathematics

OpenStudy (anonymous):

let the angles be a and b a+b =180 also a=3b+12 =>3b+12+b=180 =>4b+12=180 =>b=42 =>a=138